Accelerator Update for January 9, 2009 - January 12, 2009 |
Friday Jan 9 The day shift began with Operations monitoring the Tevatron (TeV) store 6713 (87.48E30) and the Recycler stash (300.1E10), with the Antiproton Source stacking (23.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam. At 8:51 AM, the P2 transfer line beam permit tripped off when a power supply’s (CPSF1I) output current dipped too low. Everything reset. At 9:16 AM, Operations reported that a transfer of antiprotons was lost. Operators and experts investigated. A Pbar expert reported at 10:33 AM that the problem with the lost transfer was due to an AP-1 power supply (V105). It was replaced. At 11:20 AM, a Linac Klystron RF station (KRF3) tripped off. Linac experts investigated. At 11:21 AM, the P1 transfer line beam permit tripped off when V714’s current drifted too low. Everything reset. Experts brought KRF3 back online and Linac resumed sending beam at 11:30 AM. Operations terminated store 6713 at 11:34 AM. TeV experts began a two-hour coupling study at 11:44 AM. Operations prepared Pbar for access at 12:32 PM to investigate the PLC problems. (Later, they reported finding two damaged transfer switches in the AP-10 stub area, which were repaired.) At 1:34 PM, KRF5 tripped off. It reset. Operations put the TeV into shot setup at 1:47 PM. Operations established store 6715 at 3:39 PM with an initial luminosity of 359.45E30. Operations reported at 4:01 PM that the Recycler had stashed antiprotons five times during the day shift. At 6:42 PM, a Booster Spellman power supply tripped off. Operators investigated and discovered a failed power supply. Booster experts had the power supply replaced by 8:45 PM. At 9:15 PM, Operations reported a noisy Linac quadrupole power supply (L:QPS512) that was occasionally tripping off. Operators switched its monitoring to shunt mode. Operations reported at midnight that the Recycler had stashed four times during the evening shift. Saturday Jan 10 The midnight shift began with Operations monitoring store 6715 (118.81E30) and the stash (139.5E10), with Pbar stacking (20.2E10), and with NuMI and MiniBooNE taking beam. At 12:17 AM, Operations reported that the Pelletron VME node was down. At 1:04 AM, Operations reported another noisy Linac quadrupole power supply (QPS401). Operations switched its monitoring to shunt mode. At 3:47 AM, Operations reported some Recycler sequencer problems. At 4:33 AM, KRF5 tripped off. It reset. At 5:29 AM, the P1 transfer line tripped off due to V714. Everything reset. Operations reported at 8:01 AM that the Recycler had stashed nine times during the midnight shift. At 11:40 AM, KRF1 tripped off. It reset. At 11:47 AM, KRF5 tripped off. It reset. Operations terminated store 6715 at 12:26 PM. Operations put the TeV into shot setup at 12:38 PM. A controls expert brought the Pelletron node back online at 12:41 PM. At 1:01 PM, Operations reported that beam had been falling out of MI after injection. Operators changed the vertical tune and this fixed the problem. Operations established store 6716 at 2:12 PM with an initial luminosity of 350.68E30. Operations reported at 8:05 PM that the Recycler had stashed 11 times during the last 12 hours. At 11:48 PM, LRF4 tripped off. It reset. At 11:50 PM, BRF8 tripped off. It reset. Sunday Jan 11 The midnight shift began with Operations monitoring store 6716 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 12:28 AM, KRF5 tripped off. It reset. At 3:15 AM, MIRF10 tripped off. It reset. Operations terminated store 6716 at 6:25 AM. Operations put the TeV into shot setup at 6:57 AM. Operations reported at 7:59 AM that the Recycler had stashed 12 times during the last 12 hours. Operations established store 6717 at 8:05 AM with an initial luminosity of 335.08E30. At 1:51 PM, the TeV sector C4 air compressor tripped off. Operators managed to get to the C4 service building and restart the compressor before it caused the store to abort. At 4:22 PM, the Pelletron VME node went down. Operations called in a controls expert. At 4:58 PM, MIRF10 tripped off for the first of four times in about two hours. It reset. A controls expert reported at 5:27 PM that the Pelletron problem was not with the VME, but with Accelnet, which he brought up. At 8:34 PM, MIRF10 tripped off. It reset. At 9:50 PM, a Linac quadrupole power supply (QPS111) tripped off for the first of two times in 11 minutes. Operators switched the monitor to shunt mode. At 10:35 PM, LRF4 tripped off. It reset. At 10:57 PM, LRF2 tripped off. It reset. Monday Jan 12 The midnight shift began with Operations monitoring store 6717 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. Operations terminated store 6717 at 1:21 AM. Operations put the TeV into shot setup at 1:33 AM. Operations established store 6718 at 3:03 AM with an initial luminosity of 344.78E30. Operations reported at 7:57 AM that the Recycler had stashed 11 times during the last 12 hours. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 01/12/2009 email Fermilab |
| Security, Privacy, Legal | |